Calculate the pH of a buffer solution made from 0.50 M HC2H3O2 and 0.80 M C2H3O2- that has an acid dissociation constant for HC2H3O2 of 1.5 x 10-5

1
Expert's answer
2021-03-10T06:51:55-0500

Q170302


Calculate the pH of a buffer solution made from 0.50 M HC2H3O2 and 0.80 M C2H3O2- that has an acid dissociation constant for HC2H3O2 of 1.5 x 10-5.


Solution:


HC2H3O2 is acetic acid. It is a weak acid. Its dissociation is given as


HC2H3O2 (aq) + H2O (l) ⇄  C2H3O2-  (aq) + H3O+ (aq)


Weak acid conjugate base


HC2H3O2 and C2H3O2-  will form a buffer solution. We can find the pH of this buffer solution by

using Henderson- Hasselbalch equation.


"pH = pK_a + log [\frac{conjugate\space base }{Weak\space acid }]"


"pH = pK_a + log [\frac{[A-] }{[HA] }]" ; where "pK_a = -log [K_a]"


Ka is the acid dissociation constant.


"pK_a = -log [1.5 * 10^{-5} ] = 4.824;"


substitute, pKa = 4.824, [HA] = 0.50M, and [A-] = 0.80M in the Henderson-Hasselbalch equation we have


"pH = 4.824 + log [\frac{0.80M }{0.50M }]"


"= 4.824 + log [{1.6}]"


"= 4.824 + 0.20412"


"= 5.028 ;"


which in the correct significant figure is 5.0.

Hence the pH of a buffer solution made from 0.50 M HC2H3O2 and 0.80 M C2H3O2-  is 5.0